If Qs = -20 + 10p, and Qd = 400 - 20p, what is the equilibrium price and quantity?


1
Expert's answer
2021-06-14T13:53:16-0400

Equilibrium Price ="Q_s=Q_d"


"-20 + 10p=400 - 20p"

"30p=420"

"p=14"

Equilibrium Price =14


Equilibrium Quantity ="400-(20\u00d714)=120"